Suppose that a and b belong to a commutative ring R with unity. If a is a unit of R and b^2 = 0, show that a + b is a unit of R.
1
Expert's answer
2021-08-16T17:03:38-0400

We look at this problem from two perspective.


Case 1 (b=0)

If b=0 then a+b = a. And since a is a unit of R, we can conclude that a+b=a is a unit of R.


Case 2 ("b \neq 0)"

Since a is a unit, let c be the inverse of a.

In order to show that a+b is a unit, it suffices to show that 1+bc is a unit.

Given that "b^2=0" , we have that "b^3 = 0" because "b^3=b^2b=0(b)=0"

Also, "(bc)^3 = 0"

But

"1= 1+ (bc)^3 = (1+bc)(1-bc+(bc)^2)"

That is 1+bc is a unit. Hence, a+b is a unit.